OpenStudy (anonymous):

What is an estimate of the solution of the equation 6n + 3 = 2? Use a table

OpenStudy (anonymous):

what table?

OpenStudy (anonymous):

the answer is 0.166666

jhonyy9 (jhonyy9):

6n+3=2 6n=-1 n=-1/6

OpenStudy (anonymous):

6n +3 = 2 6n = -1 n = -1/6

OpenStudy (anonymous):

Yeah, thats how u get 0.16666